Without actual divison prove that 2x2-6x2+3x2+3x-2 is exactly divisible by[x2-3x+2].
1
Expert's answer
2012-05-29T08:36:58-0400
2x2-6x2+3x2+3x-2=-x^2+3x-2=-(x^2-3x+2)& and it's obviously divisible by x2-3x+2